# The Modularity of an Abelian Variety

### v1

@@ A $2g$-dimensional $\ell$-adic Galois representation $\rho$ of $G_{\mathbb{Q}}$ given by
${\rm Tr}\left(\rho({\rm Frob}_{p})\right)=a(pI_{g};F)\quad{\rm and}\quad\det\left(\rho({\rm Frob}_{p})\right)=p^{g}\quad{\rm for\ any\ prime}\ p\not|\,\ell N,$
@@
(2.2) $[\Gamma_{g,0}(N),g+1]_{0}\cong\Omega^{\nu}\left({\mathcal{A}}_{g,0}^{\rm tor}(N)\right),\qquad\nu=\frac{g(g+1)}{2}.$
@@
for all $\gamma=\begin{pmatrix}A&B\\ C&D\end{pmatrix}\in\Gamma_{g,0}(N)$ and $\Omega\in\mathbb{H}_{g}.$ Thus $F\in[\Gamma_{g,0}(N),g+1].$ It was shown by Freitag [10] that $\omega$ can be extended to a holomorphic $\nu$-form on ${\mathcal{A}}_{g,0}^{\rm tor}(N)$ if and only if $F$ is a cusp form in $[\Gamma_{g,0}(N),g+1]_{0}.$ Indeed, the mapping
$[\Gamma_{g,0}(N),g+1]_{0}\longrightarrow\Omega^{\nu}\left({\mathcal{A}}_{g,0}^{\rm tor}(N)\right),\qquad F\mapsto F\,\omega_{0}$
is an isomorphism as complex vector spaces. We observe that if $\omega_{k}:=G(\Omega)\,\omega_{0}^{\otimes k}$ is a $\Gamma_{g,0}(N)$-invariant holomorphic form on $\mathbb{H}_{g}$ of degree $k\nu$, then $G(\Omega)\in[\Gamma_{g,0}(N),k(g+1)]_{0}$ is a cusp form of weight $k(g+1).$ Therefore according to (2.1) and (2.2), we have
@@
Let A be a simple abelian variety of dimesion $g$ defined over $\mathbb{Q}$. The following statements are equivalent : (MAV1) $A$ is modular. (MAV2) There exists a non-constant holomorphic mapping ${\mathcal{A}}_{g,0}^{\rm tor}(N)\longrightarrow A$ for some
positive integer $N$. (MAV3) There exists a non-constant holomorphic mapping $J_{g,0}(N)\longrightarrow A$ for some
positive integer $N$. (MAV4) $\rho_{A,\ell}$ is modular for any prime $\ell$.

### v2

@@ A $2g$-dimensional $\ell$-adic Galois representation $\rho$ of $G_{\mathbb{Q}}$ given by
[…]
@@
(3.3) $\Gamma_{g,0}(N),g+1]_{0}\cong\Omega^{\nu}\left({\mathcal{A}}_{g,0}^{\rm tor}(N)\right),\qquad\nu=\frac{g(g+1)}{2}.$
@@
for all $\gamma=\begin{pmatrix}A&B\\ C&D\end{pmatrix}\in\Gamma_{g,0}(N)$ and $\Omega\in\mathbb{H}_{g}.$ Thus $F\in[\Gamma_{g,0}(N),g+1].$ It was shown by Freitag [25] that $\omega$ can be extended to a holomorphic $\nu$-form on ${\mathcal{A}}_{g,0}^{\rm tor}(N)$ if and only if $F$ is a cusp form in $\Gamma_{g,0}(N),g+1]_{0}.$ Indeed, the mapping
$\Gamma_{g,0}(N),g+1]_{0}\longrightarrow\Omega^{\nu}\left({\mathcal{A}}_{g,0}^{\rm tor}(N)\right),\qquad F\longmapsto F\,\omega_{0}$
is an isomorphism as complex vector spaces. We observe that if $\omega_{k}:=G(\Omega)\,\omega_{0}^{\otimes k}$ is a $\Gamma_{g,0}(N)$-invariant holomorphic form on $\mathbb{H}_{g}$ of degree $k\nu$, then $G(\Omega)\in[\Gamma_{g,0}(N),k(g+1)]_{0}$ is a cusp form of weight $k(g+1).$
Therefore according to (3.2) and (3.3), we have
@@
Let A be a simple abelian variety of dimesion $g$ defined over $\mathbb{Q}$. The following statements are equivalent :
(MAV1) $A$ is modular.
(MAV2) There exists a non-constant holomorphic mapping ${\mathcal{A}}_{g,0}^{\rm tor}(N)\longrightarrow A$ for some
positive integer $N$.
(MAV3) There exists a non-constant holomorphic mapping $J_{g,0}(N)\longrightarrow A$ for some
positive integer $N$.
(MAV4) $\rho_{A,\ell}$ is modular for any prime $\ell$.